# On the solution of the Collatz problem

### v26

@@ Abstract.
In this paper, we first prove that given a natural number $m$ and an odd number $t$ not divisible by $3$, there exists a unique Collatz’s Sequence
@@
Then in two ways we prove that each Collatz’s Sequence always returns to 1, i.e., the Collatz problem is solved.
@@ Introduction
A problem posed by L. Collatz in 1937, also called the $13n+$ 1 problem. Let $m_{0}^{*}$ be an integer. Then one form of Collatz problem asks if iterating
(0.1) $m_{k}^{*}=\{\begin{array}[]{cc}m_{k-1}^{*}/2&\texttt{ for }m_{k-1}^{*}\texttt{even }\\ 3m_{k-1}^{*}+1&\texttt{ for }m_{k-1}^{*}\texttt{ odd }\end{array}$
@@
The members of the sequence produced by the Collatz are sometimes known as hailstone numbers. Conway proved that the original Collatz problem has no nontrivial cycles of a positive length $<400$. Lagarias (1985) showed that there are no nontrivial cycles with length $<275000$. Conway (1972) also proved that Collatz-type problems can be formally undecidable. Kurtz and Simon (2007) proved that a natural generalization of the Collatz problem is undecidable; unfortunately, this proof cannot be applied to the original Collatz problem [3, 4]. See also the work by Tao (2020) for dramatic recent progress on the problem.
@@
by which all odd numbers can be expressed. Then, we solve the Collatz problem in two ways:
(1) In the proof of Theorem 2.1, by using the matrix theory we prove that no Collatz cycle(See Definition 1.3) contains any odd number greater than 1, then we derive that each odd number $n_{i_{s}}(m_{s},t_{s})$ in a series of Collatz’s Sequences
$S_{c}(s)=\{S_{c}(m_{j},t_{j})|j=0,1,2,\ldots,s-1\}$
(See Definition 1.2) is finite and equal to 1 for $s$ finite or $s<\infty$, and finally we prove that each odd number $n_{i_{s}}(m_{s},t_{s})$ in a series of Collatz’s Sequences $S_{c}(s)$ is finite and equal to 1 as $s$ tends to infinite by using the distribution of even numbers and the law of great numbers.
(2) In the proof of Theorem 2.2, we prove that all odd numbers in the range $(N,\rho N]$(See Definition 1.7) return to 1 where lemmas 3.17-3.20 form mutual references or nested recursions to themselves, which make complexities of the Collatz problem, and finally by letting $N_{0}=N$ and $N_{i+1}=\rho N_{i}$ each odd number in the new range $(N_{i+1},\rho N_{i+1}]$ returns to 1 for $i=0,1,2,\ldots$. So we prove that all odd numbers return to 1.
@@ 1. Definitions of Collatz’s Sequences
Given a natural number $m$ and an odd number $t$ not divisible by $3$, let a Collatz’s Sequence be
@@
Given a natural number $m$ and an odd number $t$ not divisible by $3$, let a series of Collatz’s Sequences be
@@
Let a Collatz cycle be a cycle starting from an odd number $n$ and returning to $n$ by the Collatz algorithm.
@@
Let define $r_{c}(m,t,i,1)=n_{i}(m,t)/n_{i_{1}}(m_{1},t_{1})$ where $0\leq i\leq m$ and $0\leq i_{1}\leq m_{1}$, and $r_{c}(m,t,i,s)=n_{i}(m,t)/n_{i_{s}}(m_{s},t_{s})$ where $0\leq i\leq m$ and $0\leq i_{s}\leq m_{s}$.
@@
Let define $r_{c}(m_{j},t_{j})=n_{i_{j}}(m_{j},t_{j})/n_{i_{j+1}}(m_{j+1},t_{j+1})$ for $j=0,1,2,\ldots,s-1$. Then there is
@@ 2. Theorems of Collatz’s Sequences
(2.1) $m_{k}^{*}=\{\begin{array}[]{cc}m_{k-1}^{*}/2&\texttt{ for }m_{k-1}^{*}\texttt{even }\\ 3m_{k-1}^{*}+1&\texttt{ for }m_{k-1}^{*}\texttt{ odd }\end{array}$
@@
Next, by Lemma 3.9, no Collatz cycle contains any odd number greater than 1.
Then, by Lemma 3.11, given a Collatz’s Sequence $S_{c}(m,t)$, there exist a unique series of Collatz’s Sequences
$S_{c}(s)=\{S_{c}(m_{j},t_{j})|j=0,1,2,\ldots,s-1\}$
where $m_{0}=m$, $t_{0}=t$, and for $j=0,1,2,\ldots,s-1$ the odd number $n_{i_{j}}(m_{j},t_{j})$ is greater than 1, and the even number
and the odd number $n_{i_{s}}(m_{s},t_{s})$ is equal to 1 for $s$ finite or $s<\infty$.
Finally, by Lemma 3.12, each odd number $n_{i_{s}}(m_{s},t_{s})$ in a series of Collatz’s Sequences $S_{c}(s)=\{S_{c}(m_{j},t_{j})|j=0,1,2,\ldots,s-1\}$ is finite and equal to 1 as $s$ tends to infinite, and no Collatz cycle contains any odd number greater than 1 in $S_{c}(s)$, so that each initiate odd number $n_{0}(m,t)$ returns to 1.
Theorem 2.2.
Let $m_{0}^{*}$ be an integer. Then iterating
(2.2) $m_{k}^{*}=\{\begin{array}[]{cc}m_{k-1}^{*}/2&\texttt{ for }m_{k-1}^{*}\texttt{even }\\ 3m_{k-1}^{*}+1&\texttt{ for }m_{k-1}^{*}\texttt{ odd }\end{array}$
Proof.
First, given a Collatz’s Sequence
by Lemma 3.1, there are
$n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1\texttt{ for }i=0,1,2,\ldots,m,m+1,$
and
$n_{m+1}(m,t)=(3n_{m}(m,t)+1)/2=3^{m+1}t-1=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})$
By Lemma 3.3 the odd number $t$ can not be divisible by $3$. By Lemma 3.4 the odd number $n_{0}(m,t)$ can not be expressed as $16x-$ 1 nor $56x+$ 5.
First, by Lemma 3.14, each odd number $n_{0}(m,t)=4k+1$ in the range $(N,\rho N]$ returns to 1, so that half of odd numbers in the range $(N,\rho N]$ return to 1.
Second, by Lemma 3.15, each odd number $n=6x-1$ in the range $(N,\rho N]$ returns to 1, so that another $1/6$ of odd numbers in the range $(N,\rho N]$ return to 1.
@@
Third, by Lemma 3.16 all odd numbers $n_{0}(m,t)=4k+3$ in the range $(N,\rho N]$ return to 1, so that another $1/3$ of odd numbers in the range $(N,\rho N]$ return to 1.
@@ 3.1. Formation of Collatz’s Sequences
Hence, when the positive integer $m$ is finite, the number of elements of the Collatz’s Sequence $S_{c}(m,t)$ is finite and equal to $m+1$.
@@
Lemma 3.2 (Property of the natural number $m$).
@@
On the other hand, given an odd number $n_{i}(m,t)$, natural numbers $m$ and $i$, and an odd number $t$ can be determined by the formula
$n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1\texttt{ or }n_{i}(m,t)+1=2^{m+1-i}\times 3^{i}t$
@@
Since $n_{0}(m,t)=2^{m+1}t-1=2k+1$, there is an equation $12^{m}t=k+$ 1. If $k$ is even or $n_{0}(m,t)=4k+1$ then there are $m=0$ and $n_{m+1}(m,t)=2^{r}(2k_{1}+1)$ where $r$ is a positive integer, else for $k$ odd or $n_{0}(m,t)=4k+3$ there is $m>0$.
When $m>0$, there are $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1=2k+1$ and $12^{m-i}\times 3^{i}t=k+$ 1 for $i=1,2,\ldots,m$. If $i<m$ then $k$ is odd so that $n_{i}(m,t)=4k_{i}+3$, else $k$ is even so that $n_{m}(m,t)=4k_{m}+1$ and $n_{m+1}(m,t)=2^{r}(2k_{m+1}+1)$ where $r\geq 1$.
On the other hand, given an odd number $n_{i}(m,t)$, natural numbers $m$ and $i$, and an odd number $t$ can be determined by the formula
$n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1\texttt{ or }n_{i}(m,t)+1=2^{m+1-i}\times 3^{i}t$
@@
where $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m+1$, the odd number $n_{0}(m,t)\in S_{c}(m,t)$ can not be expressed as $16x-$ 1 nor $56x+$ 5.
@@
Hence the odd number $n_{0}(m,t)$ can not be expressed as $16x-$ 1 nor $56x+$ 5.
@@
where $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m+1$, let the even number $n_{m+1}(m,t)=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})$ where $r_{1}>0$, $0\leq i_{1}\leq m_{1}$, and $n_{i_{1}}(m_{1},t_{1})$ is odd and not dividable by 3.
Then a $1/2^{r}$ of positive integers $r_{1}$ satisfy $r_{1}=r$ for $r=1,2,\ldots$.
@@
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}4k+3^{m+1}-1=2(3^{m+1}2k+\sum_{j=0}^{m}3^{j})$
where $\sum_{j=0}^{m}3^{j}$ is odd, or when $m$ is odd and $t=4k+3$ where $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}4k+3^{m+2}-1=2(3^{m+1}2k+\sum_{j=0}^{m+1}3^{j})$
where $\sum_{j=0}^{m+1}3^{j}$ is odd.
Second, there is $r_{1}=2$ when $m$ is even and $t=8k-1$ where $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}8k-(3^{m+1}+1)=2^{2}(3^{m+1}2k-\sum_{j=0}^{m}(-1)^{j}3^{j})$
where $\sum_{j=0}^{m}(-1)^{j}3^{j}$ is odd, or when $m$ is odd and $t=8k-3$ where $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}8k-(3^{m+2}+1)=2^{2}(3^{m+1}2k-\sum_{j=0}^{m+1}(-1)^{j}3^{j})$
where $\sum_{j=0}^{m+1}(-1)^{j}3^{j}$ is odd.
Third, there is $r_{1}=3$ when $m=4h+1-j$ and $t=16k+3^{j}$ for $j=0,1,2,3$ where $h>0$ and $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}16k+3^{4h+2}-1=2^{3}(3^{m+1}2k-\sum_{j=0}^{2h}3^{2j})$
where $\sum_{j=0}^{2h}3^{2j}$ is odd.
Forth, there is $r_{1}=4$ when $m=8h+3-j$ and $t=32k+3^{j}$ for $j=0,1,2,\ldots,7$ where $h>0$ and $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}32k+3^{8h+4}-1=2^{4}(3^{m+1}2k-5\sum_{j=0}^{2h}3^{4j})$
where $\sum_{j=0}^{2h}3^{4j}$ is odd.
Fifth, there is $r_{1}=5$ when $m=16h+7-j$ and $t=64k+3^{j}$ for $j=0,1,2,\ldots,15$ where $h>0$ and $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}64k+3^{16h+8}-1=2^{5}(3^{m+1}2k-205\sum_{j=0}^{2h}3^{8j})$
where $\sum_{j=0}^{2h}3^{8j}$ is odd.
@@
$=2^{r}(3^{m+1}2k-\frac{3^{2^{r-2}}-1}{2^{r}}\sum_{j=0}^{2h}3^{2^{r-2}j})\texttt{ where }\sum_{j=0}^{2h}3^{2^{r-2}j}\texttt{ is odd.}$.
Hence a $1/2^{r}$ of positive integers $r_{1}$ satisfy $r_{1}=r$ for $r=1,2,\ldots$.

### v32

@@ Abstract.
In this paper, we first prove that given a nonnegative integer $m$ and an odd number $t$ not divisible by $3$, there exists a unique Collatz’s Sequence
@@
Then we discuss the Collatz problem in two ways and prove that each Collatz’s Sequence always returns to 1, i.e., the Collatz problem is solved.
@@ Introduction
A problem posed by L. Collatz in 1937, also called the $3n+1$ problem. Let $m_{0}^{*}$ be an integer. Then one form of Collatz problem asks if iterating
[…]
@@
The members of the sequence produced by the Collatz are sometimes known as hailstone numbers. Conway proved that the original Collatz problem has no nontrivial cycles of a positive length <400<400. Lagarias (1985) showed that there are no nontrivial cycles with length <275000<275000. Conway (1972) also proved that Collatz-type problems can be formally undecidable. See also the work by Tao (2020) for dramatic recent progress on the problem.
@@
by which all odd numbers can be expressed. Then we discuss and solve the Collatz problem in two ways:
(1) In the proof of Theorem 2.1, by theories of the matrixes and Diophantine Equations we prove that no nontrivial Collatz cycle(See Definition 1.3) exists in a series of Collatz’s Sequences(See Definition 1.2)
$S_{c}(s)=\{S_{c}(m_{j},t_{j})|j=0,1,2,\ldots,s-1\}\texttt{ where }0<s<\infty,$
then we derive that the odd number $n_{i_{s}}(m_{s},t_{s})$ in a series of Collatz’s Sequences $S_{c}(s)$ is equal to 1 for $s$ finite, and finally we prove that the odd number $n_{i_{s}}(m_{s},t_{s})$ in a series of Collatz’s Sequences $S_{c}(s)$ is equal to 1 as $s$ tends to infinite based on the distribution of even numbers and the law of great numbers.
(2) In the proof of Theorem 2.2, we prove that first all odd numbers in the range $(N,\rho N]$(See Definition 1.7) return to 1 where these lemmas 3.22-3.25 form mutual references or nested recursions to themselves, which make complexities of the Collatz problem, and next by letting $N_{0}=N$ and $N_{i+1}=\rho N_{i}$ each odd number in the new range $(N_{i+1},\rho N_{i+1}]$ returns to 1 for $i=0,1,2,\ldots$. So we prove that all odd numbers return to 1.
@@ 1. Definitions of Collatz’s Sequences
Given a nonnegative integer $m$ and an odd number $t$ not divisible by $3$, let a Collatz’s Sequence be
@@
Given a nonnegative integer $m$ and an odd number $t$ not divisible by $3$, let a series of Collatz’s Sequences be
@@
Let a Collatz cycle be a cycle starting from an odd number $n$ and returning to $n$ by the Collatz algorithm. If a Collatz cycle contains at least one odd number greater than 1 then it is nontrivial else it is trivial.
@@
Let define
$r_{c}(m,t,i,1)=\frac{n_{i}(m,t)}{n_{i_{1}}(m_{1},t_{1})}\texttt{ and }r_{c}(m,t,i,s)=\frac{n_{i}(m,t)}{n_{i_{s}}(m_{s},t_{s})}$
where $0\leq i\leq m$, $0\leq i_{1}\leq m_{1}$, and $0\leq i_{s}\leq m_{s}$.
@@
Let define
$r_{c}(m_{j},t_{j})=\frac{n_{i_{j}}(m_{j},t_{j})}{n_{i_{j+1}}(m_{j+1},t_{j+1})}\texttt{ for }j=0,1,2,\ldots,s-1.$
Then there is
@@
Definition 1.8.
For $\xi=1,2,\cdots,s$, let define $\eta=\eta_{s}$, $\eta_{0}=0$,
$\sigma(\zeta)=\sum_{j=1}^{\zeta}r_{j}\texttt{, }\sigma(\xi,\zeta)=\sum_{j=\xi}^{\zeta}r_{j}\texttt{, and }\eta_{\xi}=\sum_{j=0}^{\xi-1}(m_{j}-i_{j}+1)\texttt{ where }0\leq i_{j}\leq m_{j}.$
Definition 1.9.
Let define a $\eta\times\eta$ nonsingular matrix
$\left(\begin{array}[]{cccc}a_{1,1}&a_{1,2}&\cdots&a_{1,\eta}\\ a_{2,1}&a_{2,2}&\cdots&a_{2,\eta}\\ \vdots&\vdots&\cdots&\vdots\\ a_{\eta,1}&a_{\eta,2}&\cdots&a_{\eta,\eta}\\ \end{array}\right)\texttt{. Then }\mathbf{A}^{*}=\left(\begin{array}[]{cccc}A_{1,1}&A_{2,1}&\cdots&A_{\eta,1}\\ A_{1,2}&A_{2,2}&\cdots&A_{\eta,2}\\ \vdots&\vdots&\cdots&\vdots\\ A_{1,\eta}&A_{2,\eta}&\cdots&A_{\eta,\eta}\\ \end{array}\right)$
is the adjoint matrix of $\mathbf{A}$ and the determinant of $\mathbf{A}$ is $Det\mathbf{A}=3^{\eta}-2^{\eta+\sigma(s)}$ where $A_{i,j}$ is the algebraic complement of the element $a_{i,j}$ in the determinant $Det\mathbf{A}$.
@@ 2. Theorems of Collatz’s Sequences
[…]
Proof.
First, by Lemma 3.1, given a Collatz’s Sequence
there are $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m,m+1$, and
$n_{m+1}(m,t)=(3n_{m}(m,t)+1)/2=3^{m+1}t-1=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})$
Second, by Lemma 3.9, no nontrivial Collatz cycle exists in any series of Collatz’s Sequences $S_{c}(s)$ where $0<s<\infty$.
Third, by Lemma 3.17, given a Collatz’s Sequence $S_{c}(m,t)$, there exist a unique series of Collatz’s Sequences $S_{c}(s)=\{S_{c}(m_{j},t_{j})|j=0,1,2,\ldots,s-1\}$ where $m_{0}=m$, $t_{0}=t$, the odd number $n_{i_{j}}(m_{j},t_{j})$ is greater than 1, the even number
for $j=0,1,2,\ldots,s-1$, and the odd number $n_{i_{s}}(m_{s},t_{s})$ is equal to 1 for $s$ finite.
Finally, by Lemma 3.18, no nontrivial Collatz cycle exists in a series of Collatz’s Sequences $S_{c}(s)=\{S_{c}(m_{j},t_{j})|j=0,1,2,\ldots,s-1\}$ and the odd number $n_{i_{s}}(m_{s},t_{s})$ in the series of Collatz’s Sequences $S_{c}(s)$ is equal to 1 as $s$ tends to infinite, so that each initiate odd number $n_{0}(m,t)$ returns to 1.
Theorem 2.2.
Let $m_{0}^{*}$ be an integer. Then iterating
[…]
@@
By Lemma 3.3 the odd number $t$ can not be divisible by $3$. By Lemma 3.4 the odd number $n_{0}(m,t)$ can not be expressed as 6x16x-1 nor $6x+5$.
Second, by Lemma 3.19, each odd number $n_{0}(m,t)=4k+1$ in the range $(N,\rho N]$ returns to 1, so that one half of odd numbers in the range $(N,\rho N]$ return to 1.
Third, by Lemma 3.20, each odd number n=6x1n=6x-1 in the range $(N,\rho N]$ returns to 1, so that another $1/6$ of odd numbers in the range $(N,\rho N]$ return to 1.
@@
Fourth, by Lemma 3.21 all odd numbers $n_{0}(m,t)=4k+3$ in the range $(N,\rho N]$ return to 1, so that another $1/3$ of odd numbers in the range $(N,\rho N]$ return to 1.
@@ 3.1. Formation of Collatz’s Sequences
Hence, when the positive integer $m$ is finite, the number of elements of the Collatz’s Sequence $S_{c}(m,t)$ is finite and equal to $m+2$.
@@
Lemma 3.2 (Property of the nonnegative integer $m$).
@@
On the other hand, given an odd number $n_{i}(m,t)$, nonnegative integers $m$ and $i$, and an odd number $t$ can be determined by the formula
[…]
@@
Since $n_{0}(m,t)=2^{m+1}t-1=2k+1$, there is an equation $2^{m}t=k+1$. If $k$ is even or $n_{0}(m,t)=4k+1$ then there are $m=0$ and $n_{m+1}(m,t)=2^{r}(2k_{1}+1)$ where $r$ is a positive integer, else for $k$ odd or $n_{0}(m,t)=4k+3$ there is $m>0$.
When $m>0$, there are $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1=2k+1$ and $2^{m-i}\times 3^{i}t=k+1$ for $i=1,2,\ldots,m$. If $i<m$ then $k$ is odd so that $n_{i}(m,t)=4k_{i}+3$, else $k$ is even so that $n_{m}(m,t)=4k_{m}+1$ and $n_{m+1}(m,t)=2^{r}(2k_{m+1}+1)$ where $r\geq 1$.
On the other hand, given an odd number $n_{i}(m,t)$, nonnegative integers $m$ and $i$, and an odd number $t$ can be determined by the formula
[…]
@@
where $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m+1$, the odd number $n_{0}(m,t)\in S_{c}(m,t)$ can not be expressed as 6x16x-1 nor $6x+5$.
@@
Hence the odd number $n_{0}(m,t)$ can not be expressed as 6x16x-1 nor $6x+5$.
@@
where $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m+1$, let the even number
$n_{m+1}(m,t)=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})\texttt{ where }r_{1}>0\texttt{, }0\leq i_{1}\leq m_{1},$
and $n_{i_{1}}(m_{1},t_{1})$ is odd and not dividable by 3.
Then $1/2^{r}$ of positive integers $r_{1}$ satisfy $r_{1}=r$ for $r=1,2,\ldots$.
@@
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}4k+3^{m+1}-1$
$2(3^{m+1}2k+\sum_{j=0}^{m}3^{j})\texttt{ where }\sum_{j=0}^{m}3^{j}\texttt{ is odd},$
or when $m$ is odd and $t=4k+3$ where $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}4k+3^{m+2}-1$
$2(3^{m+1}2k+\sum_{j=0}^{m+1}3^{j})\texttt{ where }\sum_{j=0}^{m+1}3^{j}\texttt{ is odd}.$
Second, there is $r_{1}=2$ when $m$ is even and t=8k1t=8k-1 where $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}8k-(3^{m+1}+1)$
$2^{2}(3^{m+1}2k-\sum_{j=0}^{m}(-1)^{j}3^{j})\texttt{ where }\sum_{j=0}^{m}(-1)^{j}3^{j}\texttt{ is odd},$
or when $m$ is odd and t=8k3t=8k-3 where $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}8k-(3^{m+2}+1)$
$2^{2}(3^{m+1}2k-\sum_{j=0}^{m+1}(-1)^{j}3^{j})\texttt{ where }\sum_{j=0}^{m+1}(-1)^{j}3^{j}\texttt{ is odd}.$
Third, there is $r_{1}=3$ when m=4h+1jm=4h+1-j and $t=16k+3^{j}$ for $j=0,1,2,3$ where $h>0$ and $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}16k+3^{4h+2}-1$
$2^{3}(3^{m+1}2k-\sum_{j=0}^{2h}3^{2j})\texttt{ where }\sum_{j=0}^{2h}3^{2j}\texttt{ is odd}.$
Forth, there is $r_{1}=4$ when m=8h+3jm=8h+3-j and $t=32k+3^{j}$ for $j=0,1,2,\ldots,7$ where $h>0$ and $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}32k+3^{8h+4}-1$
$2^{4}(3^{m+1}2k-5\sum_{j=0}^{2h}3^{4j})\texttt{ where }\sum_{j=0}^{2h}3^{4j}\texttt{ is odd}.$
Fifth, there is $r_{1}=5$ when m=16h+7jm=16h+7-j and $t=64k+3^{j}$ for $j=0,1,2,\ldots,15$ where $h>0$ and $k>0$ since
$n_{m+1}(m,t)=3^{m+1}t-1=3^{m+1}64k+3^{16h+8}-1$
$2^{5}(3^{m+1}2k-205\sum_{j=0}^{2h}3^{8j})\texttt{ where }\sum_{j=0}^{2h}3^{8j}\texttt{ is odd}.$
@@
$2^{r}(3^{m+1}2k-\frac{3^{2^{r-2}}-1}{2^{r}}\sum_{j=0}^{2h}3^{2^{r-2}j})\texttt{ where }\sum_{j=0}^{2h}3^{2^{r-2}j}\texttt{ is odd.}$
Hence $1/2^{r}$ of positive integers $r_{1}$ satisfy $r_{1}=r$ for $r=1,2,\ldots$.